### v1

# FedCache 2.0: Exploiting the Potential of Distilled Data in Knowledge Cache-driven Federated Learning

Abstract
Federated Edge Learning (FEL) has emerged as a promising approach for enabling edge devices to collaboratively train machine learning models while preserving data privacy. Despite its advantages, practical FEL deployment faces significant challenges related to device constraints and device-server interactions, necessitating heterogeneous, user-adaptive model training with limited and uncertain communication. In this paper, we introduce FedCache 2.0, a novel personalized FEL architecture that simultaneously addresses these challenges. FedCache 2.0 incorporates the benefits of both dataset distillation and knowledge cache-driven federated learning by storing and organizing distilled data as knowledge in the server-side knowledge cache. Moreover, a device-centric cache sampling strategy is introduced to tailor transferred knowledge for individual devices within controlled communication bandwidth. Extensive experiments on five datasets covering image recognition, audio understanding, and mobile sensor data mining tasks demonstrate that (1) FedCache 2.0 significantly outperforms state-of-the-art methods regardless of model structures, data distributions, and modalities. (2) FedCache 2.0 can train splendid personalized on-device models with at least 28.6 improvement in communication efficiency. Our code is available at https://github.com/poppanda/FedCache2.0.
1 Introduction
Federated Edge Learning (FEL) tak2021federated; duan2023combining represents a specialized form of Federated Learning (FL) yang2019federated designed to operate at the edge of the network, which enables edge devices (clients) to jointly train machine learning models under the coordination of an edge server (server) without sharing raw data. With the growing prevalence of mobile and Internet of Things (IoT) devices coupled with increasing concerns over data privacy, FEL has empowered wide adoption of various on-device Artificial Intelligence (AI) applications, including smart transportation transportation, healthcare healthcare, and recommendation yuan2023federated; guo2021prefer. Despite its promising potential, FEL faces significant challenges in practical deployment, primarily deriving from device-side limitations and the intricate dynamics of device-server interactions in mobile edge networks. Diverse user behaviors and device capabilities necessitate the deployment of scalable and user-adaptive models tailored to heterogeneous hardware specifications and data distributions across devices tak2021federated; duan2023combining; pflsurvey; yu2021toward. These requirements call for FEL to function efficiently across a wide range of device configurations and user preferences. Additionally, communication constraints, such as limited bandwidth and uncertain connections, further complicate these interactions tak2021federated; zhang2022federated; zhu2022resilient. Common devices such as smartwatches, mobile phones, and tablets often struggle to sustain excessive wireless communication or ensure consistent connectivity. These limitations hinder the efficiency of collaborative model updates, thereby impacting the overall performance of the FEL system and the benefits it delivers to users. While a series of works have explored the trade-offs among the aforementioned constraints mills2021multi; wu2023fedict; jin2022personalized; zhu2022resilient; wang2022accelerating, they fall short of addressing all the challenges that are simultaneously prevalent in real-world edge deployments tak2021federated; zhang2022federated.
Recently, knowledge cache-driven federated learning (FedCache) wu2024federated has emerged as a game-changer in the FEL field, offering a prevailing paradigm that revolutionizes the mainstream parameters interaction protocol mcmahan2017communication; li2020federated; reddi2021adaptive; spectral; mills2021multi; jin2022personalized; lee2022preservation; wu2024federated and allowing for the joint addressing of the aforementioned challenges. By maintaining a server-side knowledge cache that stores and organizes on-device logits as transferable knowledge, FedCache facilitates communication-efficient and heterogeneous-compatible distillation-based personalized optimization without relying on unrealistic public datasets liu2022communication; wu2022communication. In addition, FedCache eliminates the requirement for multiple devices to remain online simultaneously during training, aligning with the dynamic nature of edge environments. However, the effectiveness of FedCache is limited by the breadth of information available from the remote knowledge cache due to its logits interaction design. Its applicability across various data modalities and application tasks is also restricted by the choice of a task-specific encoder used for extracting sample hashes.
To tackle these limitations, we introduce FedCache 2.0, a novel personalized FEL architecture that improves the performance of heterogeneous on-device models with efficient communication and uncertain connection tolerance. FedCache 2.0 offers a new interaction perspective between devices and the server by incorporating the benefits from both dataset distillation lei2023comprehensive and knowledge cache-driven FL wu2024fedcache, employing privacy-preserving distilled data as the knowledge cached on the server. In this setup, devices perform dataset distillation with the assistance of cached knowledge from the remote server. The distilled data is then shared with the server, ensuring the knowledge cache remains updated with the latest information. To balance system performance and communication efficiency, a device-centric cache sampling strategy is proposed for tailoring transferred knowledge for individual devices within the constraints of available communication bandwidth. The key superiorities of FedCache 2.0 compared with the original FedCache are twofold. First, FedCache 2.0 provides richer information characterization capabilities by storing and transferring distilled synthetic data rather than logits, enabling on-device models to optimize with sufficient server-side information and achieve better precision. Second, FedCache 2.0 adopts a more generalized data anonymization method, enhancing its extensibility to a broader range of data modalities and application tasks. Our proposed architecture maintains the advantages of model heterogeneity allowance, learning personalization, uncertain connection tolerance, and efficient communication from FedCache, while also achieving remarkable performance gains by fully exploiting the knowledge from distilled data.
Contributions. The main contributions of this paper are as follows: (1) We propose FedCache 2.0, a novel knowledge cache-driven federated learning architecture for personalized edge intelligence scenarios, enabling efficient post-distillation data-based knowledge storage, organization, and personalized model training while protecting data privacy. (2) We introduce federated dataset distillation and device-centric cache sampling that matches the FedCache 2.0 architecture. The former generates privacy-preserving and semantically-informed synthetic data, while the latter personalizes knowledge cache sampling to accommodate clients with diverse local data distributions and communication constraints. (3) We conduct comprehensive experiments on five datasets, encompassing image recognition, audio understanding, and mobile sensor data mining tasks. Built upon diversified data heterogeneity, model settings, and application scenarios, FedCache 2.0 not only consistently outperforms state-of-the-art methods (at least 1.7% average User model Accuracy mills2021multi enhancement) in all considered settings, but also achieves better communication efficiency (at least 29.6) compared with baseline algorithms.
2 Related Work
Personalized Federated Learning. A variety of approaches have been developed to tackle the dual challenges of learning and model scale personalization within FL. Differentiated client-side model optimization objectives are implemented in studies such as marfoq2022personalized; shi2023prior; mills2021multi; jin2022personalized; spectral; fedfed, enabling trained models to generalize across clients with varying local data distributions. Novel client-server interaction designs, which depart from the traditional FedAvg mcmahan2017communication, are explored in wu2022communication; wang2023towards; zhu2021data; wu2024exploring; huang2022learn to better accommodate diverse client hardware configurations with differently structured models. Furthermore, hybrid approaches such as wu2024fedcache; wu2023fedict are proposed to simultaneously address both model and learning personalization in FL.
Federated Learning in Edge Computing. The efficiency of executing FL at the network edge has become a hot topic. Research works such as alam2022fedrolex; ijcai2022p399; he2020group; wen2024rtifed investigate the technical frameworks required for running FL algorithms on devices constrained by computational power or storage resources. Device heterogeneity and connection uncertainty in edge environments are tackled by methodologies such as zhu2022resilient; wu2024fedcache; liu2023adaptive; wu2023fedict. Moreover, wang2022accelerating; wu2023agglomerative; deng2023hierarchical; liu2020client; wang2021resource extends FL to a multi-tier architecture involving end-edge-cloud collaborations, enhancing model training efficiency and final performance by leveraging the edge as a bridge between devices and the cloud during the training process.
Federated Learning with Alternative Information. Instead of transmitting model parameters, alternative information is utilized in the FL training process by a series of recent works. Model-agnostic outputs are exchanged between clients and the server in wu2024fedcache; itahara2021distill; huang2022learn; wu2024exploring; li2019fedmd, allowing deployment of customized on-device models across resource-heterogeneous clients. Additionally, methodologies involving uploading mixed or distilled data from clients to servers are proposed in song2023federated; oh2020mix2fld, significantly reducing communication overhead while maintaining client data privacy.
3 Problem Statement and Reformulation
Background and Preliminary. We consider an FL system deployed at the edge of the network, comprising participating edge devices (clients) coordinated by an edge server (server). Each client owns its local dataset with samples, where each sample are with data dimensions and belong to one of distinct classes. Due to differentiated user behaviors, both the local training an testing datasets among clients are non-independently and identically distributed. Throughout this paper, the terms ’device’ and ’client’ are used interchangeably. Assume that the personalized model parameters of client are denoted as , where indicates the number of parameters in the model of client . Due to system heterogeneity among devices, the required model sizes may vary across clients, such that . Each client has a local objective , which relies on its corresponding local data distribution. The overall goal is to minimize the expected objective across all clients, which is formally expressed as:
Given the instability of device connections in edge environments, multiple clients may not be online simultaneously. Besides, it is essential to minimize the communication overhead between devices and the server under the premise of guaranteeing user model accuracy mills2021multi, saving valuable wireless network resources as well as device energy.
Knowledge Cache-driven Federated Learning. We formulate knowledge cache-driven FL as a distributed optimization problem with the assistance of the remote knowledge cache on the server, that is:
where is the cross-entropy loss, is the softmax function, is the prediction function of the model on client . represents the redundant optimization component of client based on cached knowledge, with corresponding weighting term . As an example, FedCache wang2021knowledge considers model outputs (logits) as knowledge, and optimizes local models based on cached related knowledge, that is:
where is the Kullback-Leibler Divergence loss, is the -th knowledge fetched from the knowledge cache for sample index , is a hyper-parameter that controls the number of related knowledge in FedCache. However, FedCache exhibits severe limitations in providing rich, distribution-aware information for personalized optimization over devices. The amount of information attainable from the remote knowledge cache is significantly restricted due to the design of small-scale logits interactions, as shown in Figure 1 (a). This design fails to offer sufficient optimization information for clients, leading to performance bottlenecks of FedCache, as shown in Figure 1 (c). Additionally, FedCache relies on task-specific data encoders to capture private sample relations, which restricts its applicability across varied data modalities and application tasks.
FedCache 2.0 Optimization Formulation. To address the aforementioned shortcomings of FedCache, FedCache 2.0 is designed to revolutionize the transferred knowledge by shifting from logits to distilled data, as shown in Figure 1 (b). Specifically, local model optimization in FedCache 2.0 is regulated by post-sampled distilled data jointly synthesized by clients, that is:
where is the synthetic data distilled on client , represents the adaptive sample strategy tailered for client . In our design, the synthesized data after sampling serves as the knowledge that devices request from the knowledge cache. This reformulation not only provides more comprehensive semantic information for local training on clients but also enhances the control over downloaded cached knowledge, enabling task-compatible and communication-efficient personalized optimization.
4 FedCache 2.0
In this section, we introduce our proposed FedCache 2.0 with an overview illustrated in Figure 2. An execution procedure of FedCache 2.0 is elaborated in Algorithm 1.
4.1 Knowledge Cache Design
Building upon the principles of knowledge-driven FL, FedCache 2.0 caches the latest distilled data as knowledge on the server side. In terms of knowledge cache operations, we provide two operations for indexing knowledge in the cache.
Client-Based Indexing. Each client’s distilled data is indexed by their identifier, allowing for efficient updates of knowledge in the cache and prototype initialization for on-device distillation, that is:
where is the notation of the knowledge cache on the server.
Class-Based Indexing. All cached knowledge belonging to any specific class are jointly fetched, facilitating the subsequent device-centric client sampling process, that is:
where the set of all knowledge belong to class in the knowledge cache, subject to:
4.2 Federated Dataset Distillation
FedCache 2.0 introduces federated dataset distillation, which collaboratively extracts anonymous structured information from local data on individual clients. This distilled data is stored on the server for further organization and accessibility.
On-Device Dataset Distillation. All devices decompose their local models into feature extractors and classifiers. For each given sample on device , the outputs of corresponding feature extractors and classifiers are denoted as and , respectively. To start dataset distillation, each device initializes its prototype by selecting one local sample per class during the first communication round or receiving distilled data from other clients during subsequent communication rounds. The latter process is controlled by a periodically updated random replacement function , with the intermediate distilled data stored in the knowledge cache, that is:
where denotes the set of prototype samples to be optimized into synthetic data after distillation. is a subset of with elements, subject to:
Without loss of generality, we assume device sets up a prototype on class . The on-device dataset distillation process should include computing the distance between the prototype’s feature maps and those of the local data using the Gram matrix, that is:
Similarly, we compute the Gram matrix of the prototype itself:
The dataset distillation objective is then optimized following kernel ridge regression loss:
where denotes the identity matrix, and is a hyper-parameter to control the degree of regularization. Note that local data is often augmented using common dataset enhancement techniques to increase the diversity of local feature maps during distillation. After obtaining the distilled data on client , it is stored in the knowledge cache , ensuring the devices always have access to the latest distilled knowledge in the following communication rounds, that is:
In Appendix A, we will further demonstrate the privacy guarantee of distilled data transmission from devices to the server.
Collaborative Training. On-device dataset distillation relies on well-optimized feature extractors. To enhance local model performance and improve future distillation quality, devices periodically request cached distilled data from the server for personalized optimization. This collaborative training procedure is formulated as follows:
where denotes the local training loss function on client , is a gating function acting as an identity mapping when the knowledge cache is empty in the first communication round and resulting in otherwise, that is:
4.3 Device-Centric Cache Sampling
To enhance personalized performance while reducing communication overhead, we propose a device-centric cache sampling strategy that considers local data characteristics and communication budgets.
Local Label Distribution Computation. During the initialization process, each client computes its local label distribution according to its label frequency, that is:
where represents the label frequency of class on client .
Distribution-Aware Controllable Sampling. During the training process, the knowledge cache samples and distributes its stored knowledge based on , that is:
where denotes random sampling in the cached knowledge set at a probability . is a hyper-parameter ranging from 0 to 1 to control the trade-off between model performance and communication. As grows, the proportion of cached samples increases as well, leading to more cached knowledge but higher communication overhead.
Model Homo. | Method | CIFAR-10 | CINIC-10 | CIFAR-100 | |||
MTFL | 31.1 | 29.2 | 32.1 | 34.8 | 14.8 | 15.4 | |
kNN-Per | 32.7 | 34.8 | 32.8 | 29.6 | 18.8 | 18.3 | |
SCDPFL | 49.4 | 33.1 | 48.7 | 32.2 | 33.3 | 19.6 | |
FedKD | 40.9 | 23.9 | 39.2 | 22.7 | 26.1 | 14.3 | |
FedCache | 42.1 | 23.9 | 39.8 | 21.9 | 26.4 | 14.7 | |
FedCache 2.0 | 51.1 | 36.5 | 51.1 | 36.3 | 35.8 | 23.3 | |
Model Hetero. | Method | CIFAR-10 | CINIC-10 | CIFAR-100 | |||
FedKD | 39.7 | 24.1 | 39.6 | 23.6 | 26.2 | 14.1 | |
FedCache | 41.3 | 22.2 | 40.3 | 22.4 | 26.3 | 13.9 | |
FedCache 2.0 | 51.1 | 35.7 | 51.2 | 36.9 | 35.8 | 23.5 |
5 Experiments
5.1 Experimental Setup
Platform. Our experiments are conducted on a high-performance physical server equipped with 12th Gen Intel(R) Core(TM) i7-12700 CPU and multiple NVIDIA GeForce RTX 3090 GPU cards. The server’s memory consists of four 16GB Acer DDR4 modules operating at 2133 MT/s, providing a total of 64GB of RAM. Storage is handled by a KINGSTON SKC3000D2048G solid-state drive.
Datasets. We evaluate the effectiveness of our proposed FedCache 2.0 across various application tasks, including image recognition, audio understanding, and mobile sensor data mining carpineti2018custom. These experiments cover five datasets: which are CIFAR10, CIFAR100 cifar10, CINIC10 cinic10, UrbanSound8K urbansound, and TMD carpineti2018custom. Detailed illustrations of tasks and datasets are provided in Appendix B. Each complete dataset is preprocessed using the distributed data partition strategy from FedML he2020fedml, with a hyper-parameter to adjust the degree of data heterogeneity among clients.
Models. We employ five model structures, considering both deep residual network he2016deep for image data, and fully connected network for numeric data. Detailed information on model structures is illustrated in Appendix C. In addition, we consider both scenarios with homogeneous and heterogeneous on-device models in our experiments, with detailed model settings provided in Appendix D.
Baselines. We compare FedCache 2.0 against the following state-of-the-art methods: MTFL mills2021multi, KNN-Per marfoq2022personalized, spectral co-distillation for personalized FL (SCDPFL) chen2024spectral, FedKD wu2022communication and FedCache wu2024fedcache. These baseline algorithms encompass personalized/multi-task FL methods, FL algorithms addressing dual model heterogeneity and communication efficiency, and FL for edge computing. Implementation details for these algorithms are available in Appendix D.
Criteria. Following mills2021multi, we adopt the average User model Accuracy (UA) as the primary metric for evaluating model precision, focusing on the highest value achieved within 100 communication rounds. In addition, we assess communication efficiency by monitoring the learning curves, measuring average UA against per unit of communication overhead. Detailed elaborations on how communication cost is calculated are provided in Appendix E.
5.2 Performance Evaluation
Average User Model Accuracy. Table 1 displays the comparison of average UA on image recognition datasets, CIFAR10, CIFAR100, and CINIC10, with two degrees of data heterogeneity, . As displayed, FedCache 2.0 significantly outperforms all considered state-of-the-art methods across both model homogeneous and heterogeneous settings, demonstrating its superior performance and robustness in diverse edge scenarios. This substantial improvement is attributed to the enriched information characterization provided by distilled data and effective personalized optimization facilitated by device-centric cache sampling. Additional evaluations on audio understanding and mobile sensor data mining are provided in Appendix G.
Communication Cost. Figure 3 illustrates the learning curves for image recognition tasks, plotting average UA against communication cost. As shown, FedCache 2.0 exhibits significantly steeper convergence curves, reaching acceptable average UA more efficiently than competing methods, regardless of data heterogeneity, model structures, and datasets. This indicates that FedCache 2.0 can achieve robust performance improvement with reduced communication overhead, making it suitable for deployment in resource-constrained edge environments with limited wireless bandwidth. The reduction in communication cost is attributed to FedCache 2.0’s elimination of transferring cumbersome model parameters between devices and the server. Alternatively, FedCache 2.0 leverages compact distilled data as knowledge to facilitate communication-efficient personalized optimization on devices. Quantitative communication information and experimental results on additional application tasks are provided in Appendix F and Appendix G, respectively.
Method | |||||
FedCache | 41.2 | ||||
FedCache 2.0 | 51.3 | 51.7 | 51.1 | 49.8 | 48.6 |
Method | Model | CIFAR10 | CINIC10 | CIFAR100 |
FedCache | ResNet-S | 40.5 | 38.3 | 25.2 |
ResNet-S/ResNet-M/ResNet-L | 41.3 | 40.3 | 26.3 | |
FedCache 2.0 | ResNet-S | 46.6 | 46.9 | 31.5 |
ResNet-S/ResNet-M/ResNet-L | 51.1 | 51.1 | 35.8 |
5.3 Ablation Study
Impact of Cache Sampling Strategy. Table 3 presents the average UA with different values. We can conclude that increasing in the early stage generally improves the average UA due to the richer information provided by a greater number of cached samples. However, the performance gain diminishes as approaches 1. This decline is likely due to the introduction of data distribution bias across devices, which is harmful to the system’s performance.
Impact of Model Settings. Table 3 presents the average UA for different model configurations across image recognition datasets. The results indicate that heterogeneous model settings yield higher average UA compared to homogeneous settings constrained by the weakest end devices. The improvements stem from the support of more powerful devices to deploy larger and more complex models, which can make full use of computational resources among heterogeneous devices to achieve better performance. These findings underscore the benefits of model heterogeneity flexibility in FedCache 2.0.
6 Discussion
Broader Impacts. FedCache 2.0 introduces tolerance for uncertain connection, allowing devices to engage in FEL at their convenience, without relying on any two devices being online at the same time. This feature is particularly advantageous in dynamic network environments, especially for IoT applications prone to unstable connectivity from power outages or limited signal coverage. Moreover, FedCache 2.0’s generalized data anonymization broadens its utility across various data modalities and application tasks. For instance, FedCache can be seamlessly integrated into smart healthcare and e-commerce recommendation systems, facilitating personalized trained models’ deployment on smartwatches or mobile phones to monitor users’ health status and shopping preferences.
Limitations. In terms of potential limitations of FedCache 2.0, devices may maliciously upload misleading or poisoned distilled data to the server, which could negatively affect the overall system performance. In addition, the dataset distillation process conducted on devices demands considerable computational resources. This request can somewhat lead to slower training procedures on devices with low hardware capabilities or those constrained by battery life.
7 Conclusion
In this paper, we introduce FedCache 2.0, a novel personalized FEL architecture to address the challenges of resource heterogeneity, communication limitations, and dynamic network conditions in edge environments. By incorporating the benefits of both knowledge cache-driven federated learning and dataset distillation, FedCache 2.0 facilitates privacy-preserving and semantically enriched knowledge organization and transfer among devices and the server. This is achieved through an iterative process of distilling data on devices, caching them on the server, and then dispatching the cached knowledge to guide local training and subsequent distillation. Moreover, we propose a device-centric cache sampling strategy to further enhance personalized model training by adapting to client data distributions and communication constraints. Extensive experiments on various tasks and datasets demonstrate that FedCache 2.0 outperforms state-of-the-art methods with reduced communication costs, illustrating its potential as a promising solution for personalized edge intelligence scenarios.
[heading=subbibliography]
Appendix
In the appendix, we provide additional details organized as follows:
Section A illustrates the privacy guarantees of the federated dataset distillation process. It compares raw and distilled data across three image recognition datasets, highlighting how the federated dataset distillation process obscures identifiable visual information to protect user privacy.
Section B provides an overview of the tasks and datasets used in our experiments. It includes brief introductions of image recognition, audio understanding, mobile sensor data mining, and elaborate descriptions of CIFAR-10, CIFAR-100, CINIC-10, UrbanSound8K, and TMD datasets.
Section C describes the model structures used in our experiments, and provides detailed information of model parameters.
Section D provides information on the model configurations for both homogeneous and heterogeneous settings, data partition strategies, and hyper-parameter settings across the experiments.
Section E introduces the methodology for quantifying the communication costs of FL algorithms, including specific measures for different types of transmitted information like model weights, logits, sample index, and distilled data.
Section F elaborates on detailed quantitative communication costs and efficiency speed-up ratios for image recognition tasks with different degrees of data heterogeneity.
Section G provides additional evaluations of FedCache 2.0 on audio understanding and mobile sensor data mining tasks. It presents the average user model accuracy and communication efficiency for these tasks, demonstrating the effectiveness of FedCache 2.0 in diverse data modalities.
Appendix A Privacy Demonstration of Federated Dataset Distillation
In this paper, federated data distillation is employed to develop abstract semantic representations on individual devices collaboratively. Figure 4 illustrates the comparison of raw local data and distilled data on three image recognition datasets. As shown in Figure 4, raw datasets on each client feature rich and identifiable visual information, posing severe risks to user privacy. For instance, in the CIFAR-10 cifar10 dataset, the unprocessed images vividly display the characteristic features of certain common objects such as cars, birds, and plains. Similarly, in the CINIC-10 cinic10 dataset, the distinct contours of people are clearly discernible. On the contrary, the images become significantly obscured and unidentifiable after distillation. Specifically, the distilled data across all datasets exhibit a substantial reduction in distinct object contours and detailed features, preserving only rudimentary aspects of color and shape. While the transformed data preserves enough statistical attributes for effective local model training, it lacks the granularity necessary for image recovery. Hence, this transformation greatly enhances user privacy by significantly mitigating the risk of personal data leakage when such distilled data is processed centrally on servers.
Appendix B Illustration of Tasks and Datasets
In this section, we provide an overview of the tasks and datasets utilized in our experiments, which span three distinct domains: image recognition, audio understanding, and mobile sensor data mining.
Image Recognition. Image recognition is a fundamental task in computer vision, involving the assignment of images into predefined classes based on their visual content, such as color, texture, and shape. In our experiments, we evaluate our method on three image recognition datasets, which are CIFAR-10, CIFAR-100 cifar10, and CINIC-10 cinic10. All of these datasets consist of common objects in daily life, represented by vehicles, animals, and humans. They offer a diverse and challenging set of images that test the robustness and precision of FL algorithms in handling visual data.
Model | ResNet-T | ResNet-S | ResNet-M | ResNet-L | FCN-U | FCN-T |
Num. Params. | 171.0K | 265.9K | 360.8K | 455.8K | 151.3K | 162.5K |
Audio Understanding. Audio understanding involves classifying audio clips into various categories based on their acoustic properties. We employ the UrbanSound8K dataset urbansound in our experiments, which contains labeled sound excerpts from 10 distinct classes, including car horns, street music, and children playing. Each sound clip is less than four seconds long, providing a compact yet comprehensive acoustic profile for analysis. This dataset is particularly valuable for evaluating the performance of audio recognition FL systems in naturalistic urban environments.
Mobile Sensor Data Mining. Mobile sensor data mining focuses on analyzing data from mobile device sensors to infer user activities or contextual settings. In our experiments, we utilize the TMD dataset carpineti2018custom designed to identify transportation modes based on smartphone sensor readings. The TMD dataset is ideal for assessing the capability of FL algorithms to process and interpret complex, real-time sensor data in the context of mobile computing.
Appendix C Detailed Information on Model Structures
In this section, we provide detailed descriptions of the model structures used in our experiments. For experiments on image recognition, we employ deep residual network (ResNet) he2016deep with different numbers of layers to represent models with different structures, which are ResNet-T, ResNet-S, ResNet-M, and ResNet-T. For experiments on audio understanding as well as mobile sensor data mining, we adopt fully connected networks (FCN) that satisfy task-specific inputs and outputs sizes. We provide detailed information on model structures in Figure 5, and the number of parameters of adopted models in Table 4.
Appendix D Implementation Details
In this section, we outline the implementation details of our experiments, elaborating on the model configurations for both homogeneous and heterogeneous settings, the adopted data partition strategy, and the hyper-parameters employed across our experiments.
Model Configurations. We conduct experiments in both model-homogeneous and model-heterogeneous settings, respectively. For image recognition experiments with homogeneous models, all clients employ ResNet-L. For image recognition experiments with heterogeneous models, models on clients are evenly distributed among ResNet-S, ResNet-M, and ResNet-L. Note that for this set of experiments, we only consider FedKD wu2022communication and FedCache wu2024fedcache as baseline algorithms to compare with. This choice is based on the fact that MTFL mills2021multi, kNN-Per marfoq2022personalized, and SCDPFL chen2024spectral support only homogeneous models across clients. In addition, ResNet-T is adopted to facilitate the interaction of model parameters between clients and the server within the FedKD algorithm in all image recognition experiments. For audio understanding on UrbanSound8K urbansound and mobile sensor data mining on TMD carpineti2018custom, we exclusively employ the model-homogeneous setting, with all clients deploying a uniform FCN model, whose structures are denoted as FCN-U and FCN-T, respectively.
Method | Source of Code | Hyper-Parameters | Value |
MTFL | mtflurl | optimizer | customized |
batch size | 64 | ||
learning rate | 0.01 | ||
local epoch | 1 | ||
communicaion round | 100 | ||
server_lr | 0.01 | ||
bn_private | usyb | ||
0.9 | |||
0.999 | |||
0.5 | |||
kNN-Per | knnperurl | optimizer | Adam |
batch size | 64 | ||
learning rate | 0.01 | ||
local epoch | 1 | ||
communicaion round | 100 | ||
aggregator_type | centralized | ||
client_type | KNNPerClient | ||
SCDPFL | scdpflurl | optimizer | Adam |
batch size | 64 | ||
learning rate | 0.01 | ||
local epoch | 2 | ||
communicaion round | 100 | ||
0.4 | |||
0.3 | |||
FedKD | fedkdurl | optimizer | Adam |
batch size | 64 | ||
learning rate | 0.01 | ||
local epoch | 1 | ||
communicaion round | 100 | ||
tmax | 0.98 | ||
tmin | 0.95 | ||
FedCache | fedcacheurl | optimizer | Adam |
batch size | 64 | ||
learning rate | 0.01 | ||
local epoch | 1 | ||
communicaion round | 100 | ||
1.5 | |||
16 | |||
FedCache 2.0 | Our Implementation | optimizer | Adam |
batch size | 64 | ||
learning rate | 0.01 | ||
local epoch | 5 | ||
communicaion round | 15 | ||
distillation learning rate | 0.001 | ||
0.5 |
Data Partition. For all experiments conducted in this paper, we utilize the data partition strategy provided by FedML he2020fedml to simulate decentralized datasets for clients. We adopt the hyper-parameter to control the degree of data heterogeneity. A smaller number of results in a greater degree of heterogeneity among clients. Notably, the training and testing datasets for the same client have identical distributions, whereas different clients typically have different data distributions, as illustrated in Figure 6.
Hyper-parameter Settings. We provide a default setting in Table 5, which is implemented in all experiments unless stated otherwise. Note that the table only showcases a subset of hyperparameters for baseline algorithms; those not mentioned retain their original settings as specified in the corresponding open-source code. Exclusively, we uniformly set the learning rate to 0.1 and the communication round of kNN-Per to 200 for experiments on mobile sensor data mining.
Method | CIFAR10/45% | CINIC10/40% | CIFAR100/30% | |
MTFL | N/A | N/A | N/A | |
kNN-Per | N/A | N/A | N/A | |
SCDPFL | 17.4G (1.0) | 7.8G (1.0) | 17.6G (1.0) | |
FedKD | N/A | N/A | N/A | |
FedCache | N/A | N/A | N/A | |
FedCache 2.0 | 389.1M (45.8) | 193.8M (41.2) | 609.3M (29.6) | |
Method | CIFAR10/30% | CINIC10/30% | CIFAR100/15% | |
MTFL | N/A | 39.9G (1.0) | 64.5G (1.0) | |
kNN-Per | 12.1G (1.6) | N/A | 17.6G (3.7) | |
SCDPFL | 19.4G (1.0) | 10.6G (3.8) | 15.9G (4.1) | |
FedKD | N/A | N/A | N/A | |
FedCache | N/A | N/A | N/A | |
FedCache 2.0 | 315.9M (62.9) | 473.5M (86.3) | 197.5M (334.4) |
Appendix E Communication Cost Calculation
In this section, introduce how communication cost is computed in our paper. We quantify the communication cost of all considered FL algorithms in terms of pure information transmission- specifically model weights, logits, sample index, and distilled data—between clients and the server. Our communication calculation methodology leverages standard units of bytes to measure communication costs, ensuring precise and scalable metrics. For a comprehensive evaluation, we compute the communication cost of all considered algorithms where the overheads vary based on the type and frequency of transmitted information:
MTFL. Both model and optimizer parameters are serialized and transmitted in each communication round. All parameters are encoded as tensors in the float format, with each element occupying 4 bytes.
kNN-Per and SCDPFL. Similar to MTFL, these methods involve transmitting model parameters in float tensors in each communication round.
FedKD. Parameters of the student model (ResNet-T) are transmitted in each round.
FedCache. This method introduces additional complexity by transmitting sample hashes (in float format), sample index (in integer format), and logits (in float format).
FedCache 2.0. This method extends FedCache by incorporating distilled data, which is uploaded and downloaded each round. Each client uploads one synthetic data sample per class, with the volume of downloaded data regulated by the hyper-parameter . Distilled images are initially processed into a array in uint8 format (1 byte each) before conversion to JPG, optimizing communication cost.
Appendix F Quantitative Communication Information for Image Recognition
In this section, we provide detailed quantitative communication information for all considered FL algorithms in image recognition tasks with homogeneous models. We measure the communication cost in terms of the amount of information transmitted between clients and the server until the average User model Accuracy (UA) reaches a specified threshold. We also provide the communication efficiency speed-up ratio, which compares the communication cost of each method to that of the least efficient baseline method achieving the same threshold precision. Table F quantitates the corresponding efficiency of FedCache 2.0 compared with baseline algorithms. It is evident that FedCache 2.0 significantly reduces the communication cost compared to other state-of-the-art methods, with 28.6 to 332.4 communication efficiency improvements compared with state-of-the-art methods.
Method | ||
MTFL | 56.5 | 46.9 |
kNN-Per | 54.3 | 54.1 |
FedCache 2.0 | 69.4 | 64.8 |
Method | ||
MTFL | 54.8 | 46.5 |
kNN-Per | 61.6 | 48.6 |
FedCache 2.0 | 73.3 | 61.7 |
Method | /50% | /45% |
MTFL | 13.0G (1.0) | 18.4G (1.0) |
kNN-Per | 7.5G (1.73) | 7.2G (2.6) |
FedCache 2.0 | 14.3M (930.9) | 19.1M (986.5) |
Method | /50% | /45% |
MTFL | 20.5G (1.0) | 11.9G (1.0) |
kNN-Per | 12.4G (1.7) | 11.0G (1.1) |
FedCache 2.0 | 14.8M (1418.4) | 15.4M (791.3) |
Appendix G Results on Additional Application Tasks
In this section, we evaluate the effectiveness of FedCache 2.0 across additional application tasks, specifically focusing on audio understanding and mobile sensor data mining. These tasks are critical for various real-world edge applications and offer broader evaluations of FedCache 2.0 across diverse data modalities.
G.1 Evaluation on Audio Understanding
Table 8 presents the average UA on UrbanSound8K dataset under two degrees of data heterogeneity. As observed, FedCache 2.0 consistently outperforms the baseline methods, achieving higher UA across both settings. Additional experimental results on communication efficiency are provided in Table 10 and Figure 8, demonstrating significant reductions in communication cost while maintaining superior performance.
G.2 Evaluation on Mobile Sensor Data Mining
Table 8 presents the average UA on TMD dataset under two degrees of data heterogeneity. Results demonstrate that FedCache 2.0 significantly outperforms existing methods under both degrees of data heterogeneity. We also provide qualitative and quantitative evaluation of FedCache 2.0 in Figure 8 and Table 10, respectively, indicating that FedCache 2.0 not only achieves higher system performance but also requires significantly less communication overhead compared to baseline methods.
[heading=subbibliography]